We continued to enjoy good weather in the last two sessions, the focus of these two sessions was to perform Artificial Swarms on hives 2 and 7. Hive 2 has been extremely productive with the brood chamber full of brood, this colony was moved to one side and sited on stand 1, the frame with the queen and two more frames of brood were transferred to a new hive sited on the original stand. The queen took some finding, at least three trawls through the colony. The stores in two supers was split between the colonies. The two colonies were checked on 3rd June, the “swarm” has at least one capped queen cell, however no eggs were seen in the queen right colony which is a concern.
The procedure on hive 7 moved only the frame with the queen on, it took a good amount of time to find eggs on the “swarm” colony which may have chilled the brood. Examination of the swarm (now hive 6) on 3rd June revealed many perforated caps plus the queen cells were poor and of an emergency nature. Meanwhile the old queen (red) in hive 7 seems to be enjoying a new lease of life with the fresh foundation and is reported to be increasing well.
Hives 1 and 6 will now not be examined until 17th June at the earliest.
Looking at the forecast for this coming Wednesday it looks like we may well miss the showers, let’s hope so as there is much work to be done at this crucial time of the season.